# On nonlinear transitions, minimal seeds and exact solutions for the geodynamo

### v1

@@ 2.2 The optimisation procedure
$[\bm{y}^{\dagger},\bm{y}]=\int_{0}^{T_{\textrm{opt}}}\iiint(\bm{y}^{\dagger})^{T}\bm{y}\;\textrm{d}V\textrm{d}t=\int_{0}^{T_{\textrm{opt}}}\langle\bm{y}^{\dagger},\bm{y}\rangle\textrm{d}t.$ (12)
The cost functional which we wish to maximise is also split into two components as
$\mathcal{J}=\int_{0}^{T_{\textrm{opt}}}\iiint\mathcal{J}_{I}\;\textrm{d}V\textrm{d}t+\mathcal{J}_{T_{\textrm{opt}}},$ (13)
i.e. into a cumulative part defined through the time- and volume- integral of the quantity of interest $\mathcal{J}_{I}$, and another instantaneous part $\mathcal{J}_{T_{\textrm{opt}}}$ which is defined at the target time $T_{\textrm{opt}}$. In order to achieve a local maximum, all the first variations of the Lagrangian must be zero. The full details of the derivation of the adjoint equations are contained in appendix A, with the adjoint equations for our equations (3) found to be
@@
for the adjoint state $\bm{q}^{\dagger}=(\bm{u}^{\dagger},p^{\dagger},\bm{b}^{\dagger},\Pi^{\dagger},T^{\dagger})^{T}$, and where we have introduced the vorticity $\bm{\omega}=\nabla\times\bm{U}$. Note, that these equations contain a forcing term stemming from the cumulative part of the cost functional $\mathcal{J}_{I}$. The adjoint equations are solved backwards in time, with final time condition found by taking the first variation of the Lagrangian with respect to $\bm{q}(T_{\textrm{opt}})$ as
$\mathbf{M}^{T}\bm{q}^{\dagger}(T_{\textrm{opt}})=\frac{\partial\mathcal{J}_{T_{\textrm{opt}}}}{\partial\bm{q}(T_{\textrm{opt}})}.$ (15)
As $\mathcal{J}_{T_{\textrm{opt}}}=0$ for the cost functionals considered in this study the final time condition will simply be $\bm{u}^{\dagger}(T_{\textrm{opt}})=\bm{b}^{\dagger}(T_{\textrm{opt}})=\bm{0}$ and $T^{\dagger}(T_{\textrm{opt}})=0$. The boundary conditions for the adjoint variables are
$T^{\dagger}(r=r_{i})=T^{\dagger}(r=r_{o})=0,\qquad\bm{u}^{\dagger}(r=r_{i})=\bm{u}^{\dagger}(r=r_{o})=\bm{0},$ (16)
along with vacuum boundary conditions on $\bm{b}^{\dagger}$ (see §A.2). If the system has not already reached a local extremum, not all first variations of the Lagrangian will be zero. Of importance for us is the gradient of the Lagrangian with respect to the initial condition $\bm{q}_{0}$. This can be written as
$\frac{\partial\mathcal{L}}{\partial\bm{q}_{0}}=\frac{\textrm{d}\mathcal{J}}{\textrm{d}\bm{q}_{0}}=\frac{\partial\mathcal{J}_{T}}{\partial\bm{q}_{0}}+\mathbf{M}^{T}\bm{q}^{\dagger}(0),$ (17)
and gives us a way to update the initial condition to increase the value of $\mathcal{J}$, and hence move towards a local maximum.
In order to find the minimal seed that gives dynamo action, we want to find the initial condition $\bm{B}_{0}=\bm{B}(0)$ with the smallest magnetic energy that reaches a dynamo solution. In order to achieve this we maximise here different cost functionals evaluated over a time horizon $T_{\textrm{opt}}$ with respect to the initial magnetic field. Ideally, this cost functional will be maximised for seeds which yield dynamo action and will be robust to the initial guess, giving us confidence that the local maxima found via the optimisation may in fact be a global maxima (which is no way guaranteed). Details of how these different cost functionals affect the adjoint equation are detailed in appendix B. As the optimisation procedure requires that the whole direct solution be stored in memory, in order to keep the computational resources feasible we keep the time horizon $T_{\textrm{opt}}$ shorter than the time needed for a dynamo solution to be established. Therefore, following Mannix et al. (2022), we perform one long forward run after the optimisation has converged in order to determine whether the obtained initial condition results in a dynamo.
@@
Throughout the optimisation, we will require that the initial magnetic field is constrained such that $\|\bm{B}_{0}\|^{2}=M_{0}$, where $M_{0}$ is a specified magnetic energy budget. Whilst this constraint can be handled by introducing an extra Lagrange multiplier (Pringle et al., 2012), we will constrain it using Riemannian optimisation methods (Absil et al., 2007), of which rotation-based projections (Douglas et al., 1998; Foures et al., 2013) are a special case (see the discussion of Skene et al. (2022) for example). In this manner, the Euclidean gradient returned by the optimisation procedure
$\frac{\textrm{d}\mathcal{J}}{\textrm{d}\bm{B}_{0}}=\bm{b}^{\dagger}(0),$ (18)
which gives the direction of steepest ascent for $\mathcal{J}$, is projected into the Riemannian gradient that respects the constraint that the optimisation is taking place on the hypersphere $\|\bm{B}_{0}\|^{2}=M_{0}$. The minimal seed can then be found by performing a series of optimisations with decreasing values of $M_{0}$. It should be recalled though that such an approach cannot guarantee that the identified maximiser yields a global maximum as well as a local one, and it is desirable to perform multiple optimisations from different initial guesses.
@@ 2.3 Nonlinear travelling wave solutions
Many nonlinear states of interest in this system are found to take the form of travelling waves, which rotate around the $z$-axis with the non-dimensional drift frequency $\omega$ relative to our already-rotating reference frame. In order to systematically identify these states, whether they be stable or unstable, we will converge them using the Newton-hookstep algorithm (Dennis and Schnabel, 1996). To that end, we need to adopt the rotating reference frame in which the travelling waves correspond to stationary states: we thus introduce the flow map $\bm{\Phi}_{t}(\bm{q}_{0},\omega)$ which returns the state at time $t$ starting from initial condition $\bm{q}_{0}$, in a rotating reference frame with angular frequency $1+\omega$. A travelling wave solution is then an initial condition $\bm{q}_{0}$ and a drift frequency $\omega$ such that $\bm{\Phi}_{t}(\bm{q}_{0},\omega)=\bm{q}_{0}$ for all times $t$, which the Newton-hookstep algorithm identifies by computing the zeros of $\bm{f}(\bm{q}_{0},\omega)=\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)-\bm{q}_{0}$ where the time is now fixed at a small value $t_{f}$. Although the value of $t_{f}$ is arbitrary for a travelling wave (as it is stationary in the identified reference frame and so $\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)=\bm{q}_{0}$ for all $t_{f}$), a small value is used following the advice of Willis (2019a), which prevents the algorithm converging to a time-periodic solution.
@@
where $\delta$ is an (adjustable) trust region size. In this manner, the exact Newton step given by (20) is relaxed in order to constrain the size of the update to lie within the trust region. The result is a more stable algorithm that has become the primary algorithm of choice for converging unstable equilibria, travelling waves, and recurrent periodic orbits (Viswanath, 2007, 2009; Chandler and Kerswell, 2013; Budanur et al., 2017).
For our system $\bm{x}=(\bm{q}_{0},\omega)^{T}$, which means that we have one more degree of freedom, given by $\omega$, than equations $\bm{f}$. Therefore, in order to obtain a square system to solve in (21) we introduce the additional constraint that the update $\bm{\delta}\bm{x}_{i}$ cannot simply rotate $\bm{x}_{i}$ around the $z$-axis. Rotating $\bm{x}_{i}(\theta,\phi,r)$ by a small amount around $z$ axis can be achieved by setting $\bm{x}_{i}(\theta,\phi+\delta\phi,r)$, i.e. we increment $\phi$ by the small amount $\delta\phi$. Using Taylor expansions, one can write
@@ 2.4 Numerical implementation
All equations are solved using the open-source PDE solver Dedalus (Burns et al., 2020). Dedalus is a pseudo-spectral solver which is able to solve equations in spherical geometries (Vasil et al., 2019; Lecoanet et al., 2019). Dedalus parses the governing equations that have been input by the user in plain text, and provides access to the spectral discretisation and routines for timestepping them. For timestepping we use the second order semi-implicit BDF implicit-explicit multistep scheme (Wang and Ruuth, 2008). Diffusion and pressure-like terms, which allow for divergence-free conditions to be enforced, are timestepped implicitly. In this manner, divergence free conditions are solved for along with all problem variables and do not necessitate any operator splitting methods or poloidal-toroidal decompositions. Spin-weighted spherical harmonics are used for the bases in the $\theta$ and $\phi$ directions, and Jacobi polynomials are used to discretise the radial direction. The typical resolution used in our study consists of all spherical harmonics up to a maximum degree of $\ell_{\textrm{max}}=63$, Jacobi polynomial degree of $N_{\textrm{max}}=63$, and a fixed timestep of $\Delta t=0.5\times 10^{-4}$. We have verified that our code with this resolution satisfies the Boussinesq hydrodynamic and dynamo benchmark cases of Christensen et al. (2001).
@@
where $\phi$ is a scalar that is determined here by enforcing the Coulomb gauge condition $\nabla\cdot\bm{A}=0$ although other gauges may be possible (Cattaneo et al., 2020). Then the potential boundary conditions for the magnetic field on the inner and outer spheres can also be enforced directly on the vector potential (Lecoanet et al., 2019), taking the form
@@
The vector potential formulation (23) of the direct induction equations is possible since the direct equations naturally preserve the divergence of $\bm{B}$. However, the adjoint equations (14) do not naturally satisfy this property, with the divergence of the adjoint magnetic field $\bm{b}^{\dagger}$ being a gauge choice which is enforced with the Lagrange multiplier $\Pi^{\dagger}$ (see §A for more details). Hence, we solve directly for $\bm{b}^{\dagger}$. Note that $\bm{b}^{\dagger}$ must also satisfy a divergence-free constraint and continuous matching with a potential field on the spherical boundaries. A commonly used, alternative approach (Dormy, 1997; Wicht, 2002; Schaeffer et al., 2017, for example) to enforce both conditions at the same time is to decompose the considered field $\bm{X}$ into its poloidal-toroidal form as
@@
Under this decomposition, the potential boundary conditions take a very simple form in terms of the spherical harmonic decomposition of the poloidal and toroidal parts (see Hollerbach (2000) for more details). However as we do not solve the adjoint equations for the poloidal and toroidal parts since it is more natural to solve the full vector form of the equations in Dedalus, care is needed in order to properly enforce the potential boundary conditions on $\bm{b}^{\dagger}$. To that end, we make use of the relations $\nabla_{\parallel}^{2}(r\mathcal{P}^{\dagger})=\bm{e}_{r}\cdot\bm{b}^{\dagger}$, and $\nabla_{\parallel}^{2}(r\mathcal{T}^{\dagger})=\bm{e}_{r}\cdot\bm{J}^{\dagger}$, where $\nabla_{\parallel}$ is the surface Laplacian, $\bm{J}^{\dagger}=\nabla\times\bm{b}^{\dagger}$, and $\mathcal{P}^{\dagger}$ and $\mathcal{T}^{\dagger}$ are the poloidal and toroidal parts of the adjoint magnetic field, respectively. By also using the fact that the action of the surface Laplacian on spherical harmonics is $\nabla^{2}_{\parallel}Y^{m}_{\ell}=-[\ell(\ell+1)/r^{2}]Y^{m}_{\ell}$ we obtain the following boundary conditions for the adjoint magnetic field:
@@
which indirectly gives the correct conditions on the poloidal and toroidal parts of $\bm{b}^{\dagger}$.
@@
The numerical implementation of the optimisation procedure outlined in §2.2 requires a few considerations. Solving the direct equations followed by the adjoint equations provides both the value of the cost functional, and its gradient with respect to the initial magnetic field. However, it does not tell us the exact manner in which we should use the gradient to update the initial magnetic field to step towards the optimal solution. Not only must this update not change the specified energy of the initial magnetic field, but it should be done such as to increase the value of the cost functional at the next iteration. In order to handle both of these conditions we use the SphereManOpt library (Mannix et al., 2024). SphereManOpt uses techniques from Riemannian optimisation (Absil et al., 2007) in order to handle the norm constraint whilst also using line-search and optimisation algorithms in order to use the obtained gradient information to efficiently converge towards a local maximum. Based on the recommendations of Mannix et al. (2024) for continuous adjoint problems, for all optimisations performed in this study we use the conjugate gradient algorithm (Sato, 2022) with the Armijo line search routine (Wright and Nocedal, 1999).
@@
For implementing the Newton-hookstep algorithm we use the routines available in the JFNK-Hookstep Github repository (Willis, 2019b, a). Specifically, we use the FORTRAN implementation of the code developed for Openpipeflow (Willis, 2017), which is compiled into a Python module using f2py (Peterson, 2009). This enables the algorithm to be easily used together with Dedalus. The flow map $\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)$ is provided to the Newton-hookstep algorithm by using Dedalus to solve equations (3) with the Coriolis term changed to $(2+2\omega)\bm{e}_{z}\times\bm{U}$, and the boundary conditions for $\bm{U}$ now taking the form $\bm{U}(r=r_{i})=-\textit{Ek}^{-1}\omega r_{i}\sin\theta\bm{e}_{\phi}$ and $\bm{U}(r=r_{o})=-\textit{Ek}^{-1}\omega r_{o}\sin\theta\bm{e}_{\phi}$. When the Newton–hookstep algorithm converges, this corresponds to solving the equations in a reference frame that rotates with the rotation rate of the travelling wave, rather than the Earth. Hence, the fluid will appear stationary, with the boundary conditions giving the rotation rate of Earth relative to the travelling wave. Note that as with the direct equations (3), the centrifugal term is absorbed in the pressure gradient. While there are sophisticated ways for generating initial guesses for the Newton-hookstep algorithm, including those based on near recurrences (Viswanath, 2007; Chandler and Kerswell, 2013), dynamic mode decomposition (Page and Kerswell, 2020), variational methods (Lan and Cvitanović, 2004; Parker and Schneider, 2022), and convolutional autoencoders (Page et al., 2021, 2024), at the parameters used for this study we can find initial guesses directly from snapshots obtained from the evolution of the governing equations.

### v2

@@ 2.2 The adjoint-based optimal control procedure
$[\bm{y}^{\dagger},\bm{y}]=\int_{0}^{t_{\textrm{opt}}}\iiint(\bm{y}^{\dagger})^{\mathrm{T}}\bm{y}\;\textrm{d}V\textrm{d}t=\int_{0}^{t_{\textrm{opt}}}\langle\bm{y}^{\dagger},\bm{y}\rangle\textrm{d}t.$ (11)
In what follows, the cost functional we will consider is expressed as
$\mathcal{J}=\int_{0}^{t_{\textrm{opt}}}\iiint\mathcal{J}_{I}(\mathbf{q})\;\textrm{d}V\textrm{d}t,$ (12)
i.e. as the time- and volume- integral of the quantity of interest $\mathcal{J}_{I}$, with $t_{\textrm{opt}}$ a chosen time horizon for optimisation. Setting all the first variations of the Lagrangian to zero means that (i) $\bm{q}$ must evolve according to (9) (also called the “direct problem”), (ii) $\bm{q}^{\dagger}$ must solve an adjoint problem, which we will specify below for our particular system, and (iii) the initial condition for the adjoint variable $\bm{q}^{\dagger}(\bm{x},0)$ must vanish here:
$\frac{\partial\mathcal{L}}{\partial\bm{q}_{0}}=\bm{q^{\dagger}_{0}}=\mathbf{M}^{\mathrm{T}}\bm{q}^{\dagger}(\bm{x},0)=0.$ (13)
In practice, an optimum is iteratively computed by evolving the direct problem forward in time, starting from some initial (typically random) guess $\bm{q_{0}}$, then evolving the adjoint problem backward in time, down to $t=0$. At this stage, (13) provides the required gradient information to update $\bm{q_{0}}$ in order to increase the value of the cost functional $\mathcal{J}$, and hence move towards a local maximum.
With our direct problem defined by the governing MHD equations (2), and using the initial magnetic field $\bm{B}_{0}=\bm{B}(\bm{x},0)$ as the optimisation variable, we obtain the following adjoint problem (the full details of the derivation of the adjoint equations are contained in appendix A):
@@
where $\bm{q}^{\dagger}=(\bm{u}^{\dagger},p^{\dagger},\bm{b}^{\dagger},\Pi^{\dagger},T^{\dagger})^{\mathrm{T}}$ is the adjoint state, and where we have introduced the vorticity $\bm{\omega}=\nabla\times\bm{U}$. These adjoint equations, solved backward in time, are “initialised” with the final time condition (found by cancelling the first variations of the Lagrangian with respect to $\bm{q}$) that $\bm{u}^{\dagger}(t_{\textrm{opt}})=\bm{b}^{\dagger}(t_{\textrm{opt}})=\bm{0}$ and $T^{\dagger}(t_{\textrm{opt}})=0$. The boundary conditions for the adjoint variables are homogeneous Dirichlet boundary conditions for $T^{\dagger}$ and all component of $\bm{u}^{\dagger}$ on both $r=r_{o}$ and $r=r_{i}$, along with vacuum boundary conditions on $\bm{b}^{\dagger}$ and homogeneous Dirichlet boundary conditions on $\Pi^{\dagger}$ (see §A.2). Note that the integration of the adjoint equations requires that the whole direct solution be stored in memory, which can lead to high memory costs. While we did not encounter this issue within the present study, the memory requirements in future studies might become large enough to prevent the storage of the entire forward solution. This difficulty is classically overcome by the use of checkpointing schemes, which alleviate the large memory requirement at the cost of recomputing the forward solution (Griewank,, 1992). Let us mention in particular the checkpoint_schedules Python library (Dolci et al.,, 2024), which can easily be used with our developed code and has implementations of many popular optimal checkpointing schemes - including Revolve (Stumm and Walther,, 2009), disk-revolve (Aupy et al.,, 2016), periodic-disk-revolve (Aupy and Herrmann,, 2017), two-level (Pringle et al.,, 2016), H-Revolve (Herrmann and Aupy,, 2020), and mixed storage checkpointing (Maddison,, 2024).
@@
Throughout the optimisation, we will require that the initial magnetic field is constrained such that $\|\bm{B}_{0}\|^{2}=M_{0}$, where $M_{0}$ is a specified magnetic energy budget. Whilst this constraint can be handled by introducing an extra Lagrange multiplier (Pringle et al.,, 2012), we will constrain it using Riemannian optimisation methods (Absil et al.,, 2007), of which rotation-based projections (Douglas et al.,, 1998; Foures et al.,, 2013) are a special case (see the discussion of Skene et al., (2022) for example). In this manner, the Euclidean gradient returned by the optimisation procedure
$\frac{\partial\mathcal{L}}{\partial\bm{B}_{0}}=\bm{b}^{\dagger}(0),$ (15)
which provides the direction of steepest ascent for $\mathcal{J}$, is projected into the Riemannian gradient on the hypersphere $\|\bm{B}_{0}\|^{2}=M_{0}$ where optimisation is taking place.
The minimal seed that triggers dynamo action is found by performing a series of optimisations with decreasing values of $M_{0}$. The initial condition for the hydrodynamic variables are defined as an equilibrium solution of the purely hydrodynamic equations. As our time horizon is smaller than the time needed for a dynamo to be established, we perform one long forward run after each optimisation has converged (following Mannix et al., (2022)), in order to determine whether the system evolves toward a self-sustained magnetic state. The minimal dynamo seed corresponds to the optimal initial condition with smallest $M_{0}$ that still triggers a dynamo.
@@ 2.3 Nonlinear travelling wave solutions
Many nonlinear states of interest in this system are found to take the form of travelling waves, which rotate around the $z$-axis with the non-dimensional drift frequency $\omega$ relative to our already-rotating reference frame. In order to systematically identify these states, whether they be stable or unstable, we will converge them using the Newton-hookstep algorithm (Dennis and Schnabel,, 1996). To that end, we need to adopt the rotating reference frame in which the travelling waves correspond to stationary states: we thus introduce the flow map $\bm{\Phi}_{t}(\bm{q}_{0},\omega)$ which returns the state at time $t$ starting from initial condition $\bm{q}_{0}$, in a rotating reference frame with angular frequency $1+\omega$. A travelling wave solution is then an initial condition $\bm{q}_{0}$ and a drift frequency $\omega$ such that $\bm{\Phi}_{t}(\bm{q}_{0},\omega)=\bm{q}_{0}$ for all times $t$, which the Newton-hookstep algorithm identifies by computing the zeros of $\bm{f}(\bm{q}_{0},\omega)=\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)-\bm{q}_{0}$ where the time is now fixed at a small value $t_{f}$. Although the value of $t_{f}$ is arbitrary for a travelling wave (as it is stationary in the identified reference frame and so $\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)=\bm{q}_{0}$ for all $t_{f}$), a small value is used following the advice of Willis, 2019b , which prevents the algorithm converging to a time-periodic solution.
@@
where $\delta$ is an (adjustable) trust region size. In this manner, the exact Newton step given by (17) is relaxed in order to constrain the size of the update to lie within the trust region. The result is a more stable algorithm that has become the primary algorithm of choice for converging unstable equilibria, travelling waves, and relative periodic orbits (Viswanath,, 2007; Duguet et al.,, 2008; Viswanath,, 2009; Chandler and Kerswell,, 2013; Budanur et al.,, 2017).
For our system $\bm{x}=(\bm{q}_{0},\omega)^{\mathrm{T}}$, which means that we have one more degree of freedom, given by $\omega$, than equations $\bm{f}$. Therefore, in order to obtain a square system to solve in (18) we introduce the additional constraint that the update $\bm{\delta}\bm{x}_{i}$ cannot simply rotate $\bm{x}_{i}$ around the $z$-axis. Rotating $\bm{x}_{i}(\theta,\phi,r)$ by a small amount around $z$ axis can be achieved by setting $\bm{x}_{i}(\theta,\phi+\delta\phi,r)$, i.e. we increment $\phi$ by the small amount $\delta\phi$. Using Taylor expansions, one can write
@@ 2.4 Numerical implementation
All equations are solved using the open-source PDE solver Dedalus (Burns et al.,, 2020). Dedalus is a pseudo-spectral solver which is able to solve equations in spherical geometries (Vasil et al.,, 2019; Lecoanet et al.,, 2019). Dedalus parses the governing equations that have been input by the user in plain text, and provides access to the spectral discretisation and routines for timestepping them. For timestepping we use the second order semi-implicit BDF implicit-explicit multistep scheme (Wang and Ruuth,, 2008). Diffusion and pressure-like terms, which allow for divergence-free conditions to be enforced, are timestepped implicitly. In this manner, divergence free conditions are solved for along with all problem variables and do not necessitate any operator splitting methods or poloidal-toroidal decompositions. Spin-weighted spherical harmonics are used for the bases in the $\theta$ and $\phi$ directions, and Jacobi polynomials are used to discretise the radial direction. The typical resolution used in our study consists of all spherical harmonics up to a maximum degree of $\ell_{\textrm{max}}=63$, Jacobi polynomial degree of $N_{\textrm{max}}=63$, and a fixed timestep of $\Delta t=0.5\times 10^{-4}$. We have verified that our code with this resolution satisfies the Boussinesq hydrodynamic and dynamo benchmark cases of Christensen et al., (2001).
@@
where $\phi$ is a scalar that is determined here by enforcing the Coulomb gauge condition $\nabla\cdot\bm{A}=0$ although other gauges may be possible (Cattaneo et al.,, 2020). Then the potential boundary conditions for the magnetic field on the inner and outer spheres can also be enforced directly on the vector potential (Lecoanet et al.,, 2019), taking the form
@@
The vector potential formulation (20) of the direct induction equations is possible since the direct equations naturally preserve the divergence of $\bm{B}$. However, the adjoint equations (14) do not naturally satisfy this property, with the divergence of the adjoint magnetic field $\bm{b}^{\dagger}$ being a gauge choice which is enforced with the Lagrange multiplier $\Pi^{\dagger}$ (see §A for more details). Hence, we solve directly for $\bm{b}^{\dagger}$. Note that $\bm{b}^{\dagger}$ must also satisfy a divergence-free constraint and continuous matching with a potential field on the spherical boundaries. A commonly used, alternative approach (Dormy,, 1997; Wicht,, 2002; Schaeffer et al.,, 2017, for example) to enforce both conditions at the same time is to decompose the considered field $\bm{X}$ into its poloidal-toroidal form as
@@
Under this decomposition, the potential boundary conditions take a very simple form in terms of the spherical harmonic decomposition of the poloidal and toroidal parts (see Hollerbach, (2000) for more details). However as we do not solve the adjoint equations for the poloidal and toroidal parts since it is more natural to solve the full vector form of the equations in Dedalus, care is needed in order to properly enforce the potential boundary conditions on $\bm{b}^{\dagger}$. To that end, we make use of the relations $\nabla_{\parallel}^{2}(r\mathcal{P}^{\dagger})=\bm{e}_{r}\cdot\bm{b}^{\dagger}$, and $\nabla_{\parallel}^{2}(r\mathcal{T}^{\dagger})=\bm{e}_{r}\cdot\bm{J}^{\dagger}$, where $\nabla_{\parallel}$ is the surface Laplacian, $\bm{J}^{\dagger}=\nabla\times\bm{b}^{\dagger}$, and $\mathcal{P}^{\dagger}$ and $\mathcal{T}^{\dagger}$ are the poloidal and toroidal parts of the adjoint magnetic field, respectively. By also using the fact that the action of the surface Laplacian on spherical harmonic $Y_{\ell}^{m}$ is $\nabla^{2}_{\parallel}Y^{m}_{\ell}=-[\ell(\ell+1)/r^{2}]Y^{m}_{\ell}$ we obtain the following boundary conditions for the adjoint magnetic field:
@@
which indirectly gives the correct conditions on the poloidal and toroidal parts of $\bm{b}^{\dagger}$. This gives two conditions at each boundary, which, together with the boundary conditions imposed on $\Pi^{\dagger}$, gives the total number of boundary conditions needed for the adjoint equations.
@@
The numerical implementation of the optimisation procedure outlined in §2.2 requires a few considerations. Solving the direct equations followed by the adjoint equations provides both the value of the cost functional, and its gradient with respect to the initial magnetic field. This update should not change the specified energy of the initial magnetic field - which, as previously indicated, is handled by directly optimising on a Riemannian manifold. We use for this the SphereManOpt library (Mannix et al.,, 2024), similar to Pymanopt (Townsend et al.,, 2016) but focusing directly on spherical manifolds, with a conjugate gradient algorithm from (Sato,, 2022) and Armijo line search (Wright and Nocedal,, 1999).
@@
For implementing the Newton-hookstep algorithm we use the routines available in the JFNK-Hookstep Github repository (Willis, 2019a, ; Willis, 2019b, ). Specifically, we use the FORTRAN implementation of the code developed for Openpipeflow (Willis,, 2017), which is compiled into a Python module using f2py (Peterson,, 2009). This enables the algorithm to be easily used together with Dedalus. The flow map $\bm{\Phi}_{t_{f}}(\bm{q}_{0},\omega)$ is provided to the Newton-hookstep algorithm by using Dedalus to solve equations (2) with the Coriolis term changed to $(2+2\omega)\bm{e}_{z}\times\bm{U}$, and the boundary conditions for $\bm{U}$ now taking the form $\bm{U}(r=r_{i})=-\textit{Ek}^{-1}\omega r_{i}\sin\theta\bm{e}_{\phi}$ and $\bm{U}(r=r_{o})=-\textit{Ek}^{-1}\omega r_{o}\sin\theta\bm{e}_{\phi}$. When the Newton–hookstep algorithm converges, this corresponds to solving the equations in a reference frame that rotates with the rotation rate of the travelling wave, rather than the Earth. Hence, the fluid will appear stationary, with the boundary conditions giving the rotation rate of Earth relative to the travelling wave. Note that as with the direct equations (2), the centrifugal term is absorbed in the pressure gradient. While there are sophisticated ways for generating initial guesses for the Newton-hookstep algorithm, including those based on near recurrences (Viswanath,, 2007; Chandler and Kerswell,, 2013), dynamic mode decomposition (Page and Kerswell,, 2020), variational methods (Lan and Cvitanović,, 2004; Parker and Schneider,, 2022), and convolutional autoencoders (Page et al.,, 2021, 2024), at the parameters used for this study we can find initial guesses directly from snapshots obtained from the evolution of the governing equations. The travelling waves are converged to a tolerance of $10^{-6}$, with error measured using the $L_{2}$ norm of the spectral discretisation of the fields. In order to calculate the stability of each of the travelling wave solutions, we take a matrix-free approach similar to that described by Skene and Tobias, (2023). The stability problem is solved in a reference frame corotating with the travelling wave, reducing a Floquet stability problem to that of a fixed point.